(2+y)(2+y)=
also similar question
(x-5)(x-2)=

Answer by user_dude2008(1862)

(2+y)(2+y)=4+2y+2y+y^2=4+4y+y^2

(2+y)(2+y)=4+4y+y^2



(x-5)(x-2)=x^2-2x-5x+10=x^2-7x+10

(x-5)(x-2)=x^2-7x+10